OpenStudy (anonymous):

Given: <1 = (x + 7), <2 = (2x - 3),

OpenStudy (anonymous):

@Compassionate 

OpenStudy (anonymous):

can i help?

OpenStudy (anonymous):

yeah sure!

OpenStudy (anonymous):

ok. so you know that angle ABC equals x^2 right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

and you know that if you add angle 1 and angle 2, you get angle ABC

OpenStudy (anonymous):

so that means (x+7)+(2x-3)=x^2

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

now solve for x. then when you find that x, you can find x^2. then you use the x (not x^2!) to figure out what 5x-4 is. when you do those, you can see if x^2 is equal to 5x-4. if they are, then you have proven that they are congruent angles

OpenStudy (anonymous):

Ok so its 3x + 10 =x^2 So do you subtract 10 from both sides?

OpenStudy (anonymous):

be careful there! its -3 not +3.

OpenStudy (anonymous):

so 3x + 4 = x^2.

OpenStudy (anonymous):

Oh! Ok So now do I subtract 4 from both sides

OpenStudy (anonymous):

yup, you subtract the 4, but you also subtract the 3x. so you will get x^2 - 3x - 4 = 0.

OpenStudy (anonymous):

then you use the quadratic formula to solve for x. do you remember the quadratic formula?

OpenStudy (anonymous):

Why would you also substract the 3x

OpenStudy (anonymous):

subtract*

OpenStudy (anonymous):

because it has an x^2, you have to solve by using the quadratic formula

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

you remember the quadratic formula, right?

OpenStudy (anonymous):

yeah give me a second to figure the problem out

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Wait so A is 1, b is 3, and c is -4? then i plug that into the equation?

OpenStudy (anonymous):

b is -3

OpenStudy (anonymous):

yeah sorry i meant -3

OpenStudy (anonymous):

yup, you plug the a, b, c into the equation

OpenStudy (anonymous):

So x = 4, -1

OpenStudy (anonymous):

right. but if you have -1, you will end up with a negative angle, so we cant have that. So the answer is 4. So x^2 is 16, right? which means angle ABC is 16. now angle D is 5x-4. plug 4 in for that, what do you get?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

right. so if 16 = 16, you just figured it out! congratulations!

OpenStudy (anonymous):

Thanks so much for your help!

OpenStudy (anonymous):

no problem, anytime! by the way, Lady gaga is more weird than awesome :P

OpenStudy (anonymous):

Haha she has her own unique style

OpenStudy (anonymous):

lol